OpenStudy (anonymous):

Does series infinity lim n =1 of 2 +(-1)^n / n *sqrt(n) converge?

OpenStudy (anonymous):

\[\sum_{n=1}^{}\]

OpenStudy (zarkon):

is it \[\sum_{n=1}^{\infty}\frac{2+(-1)^n}{n\sqrt{n}}\]